OpenStudy (anonymous):

the length of a rectangle is three times the width. the sum of the length and the width is 24 inches. what is the lengthof the rectangle

OpenStudy (anonymous):

What rule do you need here?

OpenStudy (anonymous):

That is the first question to ask: what RULE do I use? What rule is used to solve this TYPE OF PROBLEM?

jigglypuff314 (jigglypuff314):

put what they gave you into equations. L = 3*W L + W = 24 then substitute to get 3W + W = 24 solve for W what would you get? :)

OpenStudy (anonymous):

This is a word problem that has words that you can "translate" into algebraic terms and symbols

OpenStudy (anonymous):

Specifically, this rule states to translate the following words into the following algebraic terms and symbols: "L":Length of rectangle "W":Width of rectangle "is": = "times": x "the sum of": + "and": + "what is": find